A central air-conditioning unit was purchased in 2012 for P25Million and has an expected life of 10 years and a salvage value of P5Million. What will be the book value at the end of 7 years for a) straight-line depreciation; b) SYD; and c) double-declining balance depreciation

TOPIC: DEPRECIATION

SUBJ: ENGINEERING ECONOMY
Answer #1

a) Straight-line depreciation amount = (25,000,000 - 5,000,000) / 7 = 2,000,000

The depreciation schedule is given below.

End of the year Depreciation Amount($) 2000000.00 2000000.00 2000000.00 2000000.00 2000000.00 2000000.00 2000000.00 2000000.0

Book value at the end of year 7 = $11,000,000

b) Sum of the year = n(n + 1) / 2 = 10(10 + 1) / 2 = 55

Depreciation base = 25,000,000 - 5,000,000 = $20,000,000

Year Depreciation Base Remaining life of the equipment Depreciation Amount ($) Book Value($) = Previous = Depreciation base *

Book value at the end of year 7 = $7,181,818.18

c) Double- declining depreciation rate = 2 / n = 2 / 10 = 0.2

Depreciation Amount($) End of the year = 0.20 * Previous year book value 5000000.00 4000000.00 3200000.00 2560000.00 2048000.

Book value at the end of year 7 = $5,242,880